### v1

# JMI at SemEval 2024 Task 3: Two-step approach for multimodal ECAC using in-context learning with GPT and instruction-tuned Llama models

Abstract
This paper presents our system development for SemEval-2024 Task 3: "The Competition of Multimodal Emotion Cause Analysis in Conversations". Effectively capturing emotions in human conversations requires integrating multiple modalities such as text, audio, and video. However, the complexities of these diverse modalities pose challenges for developing an efficient multimodal emotion cause analysis (ECA) system. Our proposed approach addresses these challenges by a two-step framework. We adopt two different approaches in our implementation. In Approach 1, we employ instruction-tuning with two separate Llama 2 models for emotion and cause prediction. In Approach 2, we use GPT-4V for conversation-level video description and employ in-context learning with annotated conversation using GPT 3.5. Our system wins rank 4, and system ablation experiments demonstrate that our proposed solutions achieve significant performance gains. All the experimental codes are available on Github.
[…]
1 Introduction
Emotion Cause Analysis (ECA) is centered around the extraction of potential cause clauses or pairs of emotion clauses and cause clauses from human communication, enabling a deeper understanding of communication dynamics. By incorporating multimodal cues like visual scenes, facial expressions, and vocal intonation, it facilitates a comprehensive and technically robust analysis of the factors that trigger diverse emotional reactions (Mittal et al., 2021; Zhang and Li, 2023; Zheng et al., 2023b). Despite the considerable amount of research conducted using diverse audio, visual, and text modalities (Gui et al., 2018; Xia and Ding, 2019; Fan et al., 2020; Shoumy et al., 2020; Abdullah et al., 2021), there has been a noticeable gap in the exploration of multimodal ECA in natural settings (human conversations). In this context, Wang et al. (2023a) introduce Multimodal Emotion Cause Analysis in Conversations (ECAC) task and provide Emotion-Cause-in-Friends (ECF) dataset, which incorporates text, audio, and video modalities. This task consists of two sub-tasks: Textual Emotion-Cause Pair Extraction in Conversations (Subtask 1) and Multimodal Emotion Cause Analysis in Conversations (Subtask 2). A detailed description of these sub-tasks can be found in the task description paper (Wang et al., 2023a).
In our submission to Subtask 2 of multimodal ECAC, this paper presents two distinct approaches to address the ECAC problem, giving competitive results. Drawing inspiration from the effectiveness of LLMs in diverse downstream tasks (Wang et al., 2023b, 2024; Yang et al., 2024), including emotion recognition, we propose two LLM-based approaches that decompose the emotion-cause pair extraction process into two steps. The first step involves predicting the emotions of the utterances in the conversation. In the next step, we utilize these emotion labels to guide cause extraction. Approach 1 involves instruction-tuning two separate Llama 2 models for emotion and cause prediction, while Approach 2 leverages the in-context learning (ICL) capabilities (Dong et al., 2023) of the GPT-3.5 model. Additionally, we introduce an efficient technique using the GPT-4V model to extract conversation-level descriptions from video modality.
During the evaluation, our team ranked 4th on the leaderboard competing against more than 25 teams with a weighted-F1 score of 0.2816.
2 Background
Task definition
The input for the task, $D$, comprises N conversations. As described by Wang et al. (2021), given a conversation $D_{i}=\{u_{1},u_{2},\dots,u_{M}\}$ consisting of $M$ utterances, where each utterance is represented by text, audio, and video, i.e. $u_{j}=[t_{j},a_{j},v_{j}]$, the goal of the task is to extract a set of emotion-cause pairs $P=\{\dots,(u_{k}^{e},u_{k}^{c}),\dots\}$, where $u_{k}^{e}$ denotes an emotion utterance and $u_{k}^{c}$ corresponds to the cause utterance.
Dataset
We use the Emotion-Cause-in-Friends (ECF) dataset provided by Wang et al. (2023a), which is summarized in Table 3 in Appendix B. This dataset contains 13,509 multimodal utterances that occur in the American sitcom Friends with 9272 emotion-cause pairs. Each utterance consists of the text, video, and audio.
3 Methodology
3.1 Overview
We treat the task at hand as a two-step process. In the first step, we predict the emotion of each utterance in all N conversations. Here, the context $C_{j}$ for utterance $u_{j}$ of conversation $D_{i}$ is the entire conversation itself. Given $E$ target emotion labels and $\hat{y_{j}^{e}}$ as the predicted emotion label, the problem can be formulated as:
$\hat{y_{j}^{e}}=\arg\max_{e}\mathcal{P}(y^{e}|u_{j},C_{j},\theta)$ (1)
where $\theta$ denotes the parameters.
In the second step, given these emotion labels, we predict the causes of each utterance that has an emotion other than neutral. The causes will be a subset of all utterances in the conversation $D_{i}$. Let the learned function be $f:U\rightarrow 2^{U}$, where $U$ is the set of all utterances in the given conversation. It predicts the subset $\hat{y_{j}^{c}}$ of cause of emotion utterance $u_{j}$ as:
$\hat{y_{j}^{c}}=\arg\max_{y^{c}\in 2^{U}}\mathcal{P}(y^{c}|u_{j},\hat{y_{j}^{e}},C_{j},\theta)$ (2)
where $\hat{y_{j}^{e}}\neq$ neutral.
3.2 Approach 1: Fine-tuned Llama-2
In our first approach, we perform instruction fine-tuning of the Llama 2 Large Language Model, an open-source model developed by GenAI, Meta (Touvron et al., 2023). From the three variants with 7, 13, and 70 billion parameters, we use the 13 billion parameter model due to resource constraints, albeit the performance of this model achieves state-of-the-art results on various downstream NLP tasks compared to other models of similar sizes. In addition, we use the Llama 2-chat version of the model [^1] https://huggingface.co/meta-llama/Llama-2-13b-chat-hf, which is optimized for dialogue use cases as it aligns with our task. In our approach, we use Llama2 API [^2] https://www.llama2.ai/ for prompt engineering. Through zero-shot prompting, we select optimal prompts for emotion identification and cause prediction. We observed that treating these two tasks separately resulted in better model output. This approach involves first identifying the emotions of all utterances in the conversation. We then add these emotion labels to the conversation and prompt the model to predict the causes for each emotion utterance. Consequently, we perform supervised fine-tuning of two separate Llama 2 models for these tasks. Although this increases the inference time, the significant performance gains outweigh the introduced latency. We treat both tasks as conditional generation, where the model generated the emotion label in the first case and the cause list in the second case, given the prompt. Detailed explanations of these approaches are provided in the following sections. The fine-tuning procedure is shown in Fig.1.
Figure 1: Pipeline for fine-tuning Llama (Approach 1)
3.2.1 Emotion recognition
To perform emotion recognition, we create a dataset where each sample includes an utterance $u_{j}$ from one of the $N$ conversations $D$ for which the LLM needs to output the emotion label. We incorporate the entire conversation $D_{i}$ along with speaker information as context in our prompt. This contextual information enhances the model’s understanding of the flow of emotions within the conversation, as demonstrated by our ablation studies in Section 5. The instruction $I_{j}^{e}$, which gave the best results, is given in Appendix  F.1 along with detailed prompt examples. The prompt consists of the instruction $I_{j}^{e}$ and the context $C_{j}$ for utterance $u_{j}$:
$Prompt_{j}=(C_{j},I_{j}^{e})$ (3)
Using this prompt as the input and the corresponding true emotion label $y_{j}^{e}$, we perform supervised fine-tuning of a Llama 2-13b model.
$\hat{y_{j}^{e}}=\textbf{Llama}_{e}(Prompt_{j},\theta)$ (4)
We use a quantized version of the model due to memory limitations and perform Quantized Low-Rank Adaptation (QLoRA) (Dettmers et al., 2024) as a parameter-efficient fine-tuning technique. The training details are provided in the Section 4.
3.2.2 Cause prediction
To prepare the dataset for cause prediction, we incorporate the emotion labels obtained for each utterance. The conversation contect now includes the emotion labels for each utterance $u_{j}$ excluding those with a predicted emotion label $\hat{y_{j}^{e}}$ of neutral. This approach enhances the model’s ability to analyze causal dependencies and identify which utterances may have contributed to a specific emotion. The output for cause prediction is a list of cause utterance IDs. The instruction is provided in Appendix  F.1. The modified prompt for this step consists of this instruction $I_{j}^{c}$ along with the conversational context containing emotion labels $C_{j}^{e}$:
$Prompt_{j}=(C_{j}^{e},I_{j}^{c})$ (5)
Next, we perform supervised fine-tuning of a new Llama 2-13b model using this prompt as the input and the corresponding true list of causes:
$\hat{y_{j}^{c}}=\textbf{Llama}_{c}(Prompt_{j},\theta)$ (6)
3.2.3 Adding video captions
To integrate cues from the videos corresponding to each utterance, we experimented using video captions generated using GPT-4 Vision as additional context for the model. However, we observed a notable decrease in performance since descriptions for individual utterances were somewhat noisy and did not effectively guide the predictions. Moreover, the captions often contained multiple emotions causing confusion for the model. As a result, we do not utilize these during training.
3.3 Approach 2: In-Context-Learning GPT
Figure 2: Pipeline of In-Context-Learning GPT Method (Approach 2)
Our second approach (Fig. 2) tackles subtask 2 by obtaining conversation-level video captions using the GPT-4V(ision) model by OpenAI Yang et al. (2023). For emotion prediction, we retrieve a semantically similar conversation from the training set whose emotion annotations are explained as demonstration examples in the prompt for the GPT-3.5 model [^3] https://platform.openai.com/docs/models/gpt-3-5-turbo. For each predicted emotional utterance, we perform cause prediction within a context window around the emotional utterance. Due to the complex nature of the task, we leverage in-context-learning Dong et al. (2023) by retrieving similar context windows from the training set whose cause annotations are explained as demonstration examples in the prompt for the GPT-3.5 model. We discuss each step in the subsequent sections.
3.3.1 Video Captioning
GPT-4V has the capability to process video sequences (Yang et al., 2023; Lin et al., 2023). In our approach, we extract conversation-level captions from the videos. However, due to rate limits and the costs considerations, we use a compact image representation for each video associated with the utterances of a conversation. Therefore, these image sequences serve as input to the GPT-4V model, generating a description for the entire conversation. The prompt is shown in the Fig. 3.
Figure 3: Video Captioning Pipeline
For an utterance, we sample nine equidistant frames across its video length. These frames aim to capture the dynamics of the whole video. We arrange these frames in a $3\times 3$ grid, following a row-major order. Additionally, we include the speaker text below the grid to provide further context to GPT-4V. The process is illustrated in Fig. 3.
To accommodate the rate limits of the Vision API, we batch the utterances of a conversation and obtain outputs independently from the Vision model. We stitch all the outputs of a batched conversation into a single caption using GPT-3.5 (Appendix Fig. 16).
3.3.2 Emotion Recognition
GPT tends to be uncontrollable when performing zero-shot recognition of emotions in conversations Qin et al. (2023) outputting emotions that are not a valid category of labels. To guide and control the process, we leverage in-context learning (ICL) by retrieving a conversation from the training set whose emotions are already annotated. The emotions in this conversations are explained by GPT-3.5 (Appendix Fig. 17). This retrieved conversation and its explanation serve as a demonstration for GPT to learn from, enabling it to recognize emotions in conversations more accurately. In addition, the prompt template includes the video caption as part of the input, as shown in Appendix Fig. 18.
To ensure effective ICL, it is important to provide general and descriptive examples that aid in solving the current task. In our approach, we sampled conversations from the training set containing all emotion categories. These conversations were stored as text-embedding-ada-002 embeddings (Neelakantan et al., 2022) in a vector database. At test time, we compute the embedding for a conversation and retrieve the closest matching embedding from the database based on Euclidean distance. The retrieved embedding aids ICL in improving emotion understanding and recognition.
3.3.3 Cause Prediction
Following prediction of emotions, we proceed to predict the causes for each emotional utterance within a context window around that utterance. The bounds of the context window are discussed in Appendix C.2 We chose not to predict causes across entire conversations to save costs on tokens.
For predicting causes of an utterance with emotion $e$ within a given context window $c$, we retrieve context windows containing utterances with the same emotion $e$ that exhibit semantic similarity to $c$. This retrieval is accomplished through the Euclidean distance comparison of text-embedding-ada-002 embeddings derived from the training data. The retrieved conversation’s causes are explained by GPT-3.5 (Appendix Fig. 19). Learning from the explained retrieved-context windows, cause prediction on $c$ can be performed by GPT-3.5. Video captions are also included in the prompt (Appendix Fig. 20), since the local window may have lost some broader context.
4 Experimental setup
Training details
For approach 1, the data is split into train, test, and validation sets in the ratio 8:1:1. We use peft library  [^4] https://huggingface.co/docs/peft/en/index for Parameter-Efficient Fine-Tuning. Due to memory constraints, we fine-tune a 4-bit quantized Llama-2 model using bitsandbytes library  [^5] https://github.com/TimDettmers/bitsandbytes. We report the details of the implementation for both approaches in Appendix  C.
Evaluation metrics
For evaluating, we report the precision, recall, F1-score, and weighted F1 which can be found on the competition website. [^6] https://nustm.github.io/SemEval-2024_ECAC/
5 Results and Discussion
Main results
Both of our approaches gave competitive rankings on the official leaderboard for subtask 2 as shown in Table 1. In-context-learning GPT gave better results on the evaluation set compared to Fine-tuned Llama, thus our final position on the leaderboard was rank 4.
System w-avg F1 F1
1. Samsung Research China-Beijing 0.3774 0.3870
2. NUS-Emo 0.3460 0.3517
3. SZTU-MIPS 0.3435 0.3434
4. GPT-ICL (Ours) 0.2758 0.2816
5. MotoMoto 0.2584 0.2595
6. Fine-tuned Llama (Ours) 0.2558 0.2630
Table 1: Leaderboard Results on Evaluation Data
Ablation study
We conduct extensive ablation studies to measure the importance of the techniques we employ summarized in Table  2. For these experiments, we use a subset of our test set from the training set containing 528 utterances. It can be seen that the performance of zero-shot Llama as well as GPT is the lowest. Instruction-tuning and ICL clearly improve the performance on the task, showcasing the significance of making LLMs context-aware when tackling downstream tasks. Adding self-causes improves performance in both zero-shot and context-aware cases highlighting their importance. The detailed table is in Appendix  D.
Approach F1 w-avg F1
Zero-shot Llama
    - w/o self-causes 0.117 0.116
    - w/ self-causes 0.222 0.215
Instruction-tuned Llama
    - w/o self-causes 0.325 0.318
    - w/ self-causes 0.364 0.352
Zero-shot GPT
    - w/o self-causes 0.100 0.097
    - w/ self-causes 0.189 0.184
In-context-learning GPT
    - w/o self-causes 0.235 0.241
    -w/ self-causes 0.329 0.334
Table 2: Results on Validation Set.
Limitations
Our approaches are specific to one dataset and may not generalize well to other datasets. Due to resource limitations, we fine-tune a Llama 13b parameter model instead of 70b and use QLoRA instead of updating all parameters. To save costs, we used GPT-3.5 model instead of GPT-4. Even with extensive prompt engineering, GPT models tend to hallucinate or give unstructured outputs, requiring retry repeatedly.
6 Conclusion
We tackled the Multimodal ECAC task with a two-step framework of recognizing emotions first and then predicting their causes using LLMs. We implemented two approaches: a Llama-2 model which has been fine-tuned with instructions and a GPT model which solves the task by learning from demonstration examples in context. Conversation-level video captions were extracted to provide more context to LLMs. Our second approach was our best submission for the task, placing us at rank 4 with our first approach being placed at rank 6. Our results were under cost constraints and further investigation with larger Llama-2 models and GPT-4 with more sophisticated ICL approaches are a clear follow-up of our work.
Appendix A Related Work
Our system is designed to prioritize Subtask 2 which is directly related to text-based and multimodal ECA. In the following sections, we will present relevant research that addresses both unimodal (text-based) and multimodal ECA.
Text-based ECA
Advancements in text-based ECA (Xia and Ding, 2019; Hsu et al., 2018; Peng et al., 2022; Pereira et al., 2022) have made significant strides within the field of sentiment analysis. The task on emotion cause extraction (ECE) was initially proposed by Chen et al. (2010) on a Chinese corpus. Several studies (Li and Xu, 2014; Ghazi et al., 2015; Yada et al., 2017) have explored ECE task, using both rule-based and machine learning approaches that operate at the phrase or word level of the text data. Furthermore, (Gui et al., 2018) reformulated the ECE task as a clause-level classification problem and constructed a Chinese emotion-cause corpus based on the news data. Considering the effectiveness of clause-level units in indicating emotions, Xia and Ding (2019) introduced the task of Emotion-Cause Pair Extraction (ECPE) for extracting potential emotion-cause pairs from texts. Numerous deep learning models (Zhong et al., 2019; Wei et al., 2020; Chen et al., 2020; Singh et al., 2021; Li et al., 2021; Wang et al., 2023c) have been developed to address ECPE tasks. Additionally, graph-based approaches (Zheng et al., 2023a; Hu et al., 2021; Zhao et al., 2023) that utilize graphs to model dialog context and capture interactions between speakers and utterances hold significant potential. The focus on transformer models and the rapid progress in LLMs such as ChatGPT [^7] https://chat.openai.com/ and Llama (Touvron et al., 2023), have significantly boosted the performance of various NLP tasks (Imran et al., 2023) including ECPE (Wang et al., 2023d; Imran et al., 2023; Wu et al., 2024; Zheng et al., 2022).
Multimodal ECA
Given the strong association between facial cues and emotion, integrating modalities to improve emotion recognition has attracted a lot of attention (Sebe et al., 2005; Li et al., 2023; Zhang et al., 2023; Fu et al., 2023). Several key multimodal datasets (Wöllmer et al., 2013; Zadeh et al., 2016; Chou et al., 2017; Barros et al., 2018; Poria et al., 2019; Yu et al., 2020) have emerged to support and advance research. The availability of open conversation data has facilitated the expansion of multimodal conversation datasets, which includes various types of conversations such as dyadic interactions (Busso et al., 2008), and multi-participant communications (Hsu et al., 2018; Poria et al., 2019; Firdaus et al., 2020; Zheng et al., 2023b).
Large Language Models
The emergence of Large Language Models such as GPT-4 Achiam et al. (2023), Llama Touvron et al. (2023), PaLM Anil et al. (2023), etc. has transformed the research landscape. Recently, there has been a surge in the application of LLMs to a multitude of domains. Zhang et al. (2024) extend their capabilities to the task of emotion recognition where they fine-tune a Llama 2-7 billion parameter model for emotion prediction. Lei et al. (2023) introduce a retrieval template module along with speaker identification and emotion-impact prediction tasks to improve the performance of LLM. In our work, as part of approach 1, we develop two distinct LLM-based experts separately for emotion and cause prediction.
Qin et al. (2023) investigated the task of zero-shot emotion cause prediction using ChatGPT with limited success. Recently, a new paradigm of in-context learning (ICL) Dong et al. (2023) has emerged for LLMs that involves learning from a few examples to solve a variety of complex reasoning tasks Wei et al. (2022b), Wei et al. (2022a). Wu et al. (2024) proposed a Chain of Thought (CoT) Wei et al. (2022b) approach for emotion cause pair extraction. Our approach 2 extends the idea of ICL towards solving the task of multimodal emotion cause pair extraction in two steps.
Appendix B Dataset
We use the Emotion-Cause-in-Friends (ECF) dataset curated by Wang et al. (2023a) summarized in Table 3. This dataset contains a total of 13,509 multimodal utterances that occur in the American sitcom Friends with 9272 labeled emotion-cause pairs. Each utterance consists of the text with corresponding video and audio.
Class-distribution
The dataset is imbalanced as shown in Fig. 4 wherein around 44% of the utterances have neutral emotion. Disgust and Fear constitute only 3% and 2.7% of the emotions.
Figure 4: Percentage of each of the seven emotion categories
Relative positions of emotion and causes
Interestingly, 49.95% of the causes are self-causes meaning that the same utterance caused itself as shown in Fig. 5. This is also intuitive, as what one speaks or expresses often elicits the emotion of their utterance. Note that the dataset curators have also annotated utterances coming after the emotion utterance as its cause. These constitute only about 2.8% of all causes and are one or two utterances away. 94.95% of the causes are 0-5 utterances behind the emotion utterance. The fact that what you speak or other interlocutors in the conversation speak affects the emotion of subsequent utterances explains this phenomenon.
Figure 5: Relative position of emotion and causes
Items Number
Conversations 1344
Utterances 13,509
Emotional Utterances 7,690
Self-Causal Utterances 4,892
Non-Self-Causal Utterances 2,189
No Cause Emotional Utterances 609
Later-Causal Utterances 177
Table 3: Statistics of causes for emotional utterances.
Appendix C Implementation details
C.1 Training details for Llama
Both emotion and cause prediction training used one Nvidia A100 40GB GPU for training (Available on Google Colab Pro priced at \$11.8/month). We train for one epoch due to constraints on Colab usage with gradient accumulation steps as 8 with an effective batch size of 8. A cosine learning rate scheduler and Adam optimizer are used. Inference is performed using two Tesla T4 16GB GPUs (Available on Kaggle for free (30 hrs/month)).
The long context length of 4096 tokens of the Llama 2 models, allows us to include the entire conversation as context and input that to the model. We perform experiments to analyze the maximum token counts in the dataset and observe that they do not exceed 1600 as shown in Figure 6. In case the token count exceeds the limit for the LLM we can use a window of utterances around the given utterance as context for predicting its emotion.
Hyperparameter Value
Lora alpha 16
Lora dropout 0.1
Attention heads 16
Learning rate 1e-3
Epochs 1
LR scheduler cosine
Warmup ratio 0.03
Weight decay 0.001
Table 4: Hyperparameters for fine-tuning
Figure 6: Distribution of token counts for Llama tokenizer
C.2 Details for in-context learning GPT
We use the LangChain [^8] https://github.com/langchain-ai/langchain library to implement our three pipelines: video captioning, emotion recognition, and cause prediction. We use the interface provided by LangChain to communicate with OpenAI’s API models detailed in Table 5.
Model API Name
GPT-4V gpt-4-vision-preview
GPT-3.5 gpt-3.5-turbo-1106
Embeddings text-embedding-ada-002
Table 5: OpenAI API Model Names
Context windows
The bounds of the context window are given in Table 6. The bounds were informed by the distribution of the majority of relative positions of causes in the training set Appendix B).
Position Previous Next
Beginning 0 2
End 5 0
Middle 5 2
Table 6: Context Window Bounds in each Direction
Vector databases
For creating vector databases, we use the FAISS Library Douze et al. (2024). We created a FAISS index containing embeddings of 12 conversations from the training set which contains all emotion categories. For cause prediction, we created a FAISS index for each of the 6 emotion categories and 3 possible positions of emotional utterance giving us a total of 18 indices. Each of these indices contained embeddings of context windows (bounds defined in Table 6) from the training set corresponding to each emotion and position.
Approach P R F1 w-P w-R w-avg F1
Zero-shot Llama w/o self-causes 0.089 0.168 0.117 0.090 0.168 0.116
Zero-shot Llama w/ self-causes 0.157 0.372 0.222 0.152 0.372 0.215
Instruction-tuned Llama w/o self-causes 0.351 0.304 0.325 0.335 0.304 0.318
Instruction-tuned Llama w/ self-causes 0.360 0.367 0.364 0.342 0.367 0.352
Zero-shot GPT w/o self-causes 0.081 0.130 0.100 0.087 0.130 0.097
Zero-shot GPT w/ self-causes 0.140 0.290 0.189 0.149 0.290 0.184
In-context-learning GPT w/o self-causes 0.216 0.256 0.235 0.241 0.256 0.241
In-context-learning GPT w/ self-causes 0.261 0.445 0.329 0.280 0.445 0.334
Table 7: Results on Validation Set. P: precision, R: recall, w: weighted.
C.3 Post-Processing
In both our approaches, after getting the causes, we perform a post-processing step where we add the emotional utterance as its own cause which we call self-causes. This gives significant performance boosts as a majority of the causes are self-causes as pointed out in Appendix  B.
Appendix D Detailed results
The detailed results on precision, recall, and F1-scores are given in Table  7.
Appendix E Error Analysis
We conduct error analysis for the output of emotion recognition using the two approaches. The performance of zero-shot Llama is extremely poor where the model predicts the label joy for almost all utterances (Fig. 7). On adding the conversational context, the model can identify the emotional nuances better, yet often predicts joy or surprise for neutral (Fig. 8). Instruction fine-tuning significantly boosts performance where the model can now differentiate distinct emotions (Fig. 9). The performance on disgust and fear is low due to the class-imbalance problem. In our test subset, the support of disgust and fear is only 13, as shown in Table  8. We observed similar trends in the case of our second approach. Zero-shot GPT (Fig. 10) tends to only identify the neutral utterances accurately and fails in other categories. The incorporation of in-context learning (Fig. 11) improves the accuracy in identifying different emotion categories but there is little to no improvement in identifying disgust or anger utterances.
Figure 7: Confusion matrix for zero-shot emotion recognition without context using Llama
Figure 8: Confusion matrix for zero-shot emotion recognition with context using Llama
Figure 9: Confusion matrix for emotion recognition with context using fine-tuned Llama
Figure 10: Confusion matrix for emotion recognition using Zero-shot GPT-3.5
Figure 11: Confusion matrix for emotion recognition using GPT-ICL
Approach Metric Anger Disgust Fear Joy Sadness Surprise Neutral
Supp 71 13 13 101 54 77 209
Zero-shot Llama w/o context P 0.0000 0.0000 0.0000 0.1881 0.0000 0.0000 0.0000
R 0.0000 0.0000 0.0000 1.0000 0.0000 0.0000 0.0000
F1 0.0000 0.0000 0.0000 0.3166 0.0000 0.0000 0.0000
Zero-shot Llama with context P 0.7500 0.2857 0.0000 0.3798 0.4583 0.3200 0.5663
R 0.0845 0.1538 0.0000 0.7822 0.2037 0.5195 0.4498
F1 0.1519 0.2000 0.0000 0.5113 0.2821 0.3960 0.5013
Fine-tuned Llama with context P 0.5641 0.0 0.3333 0.6210 0.625 0.6103 0.6666
R 0.6197 0.0 0.1538 0.5842 0.3704 0.6104 0.7943
F1 0.5906 0.0 0.2105 0.6020 0.4651 0.6104 0.7249
Zero-Shot GPT P 0.5652 0.2500 0.2727 0.4265 0.5385 0.5200 0.5906
R 0.3333 0.4000 0.4286 0.5370 0.1842 0.3023 0.7426
F1 0.4194 0.3077 0.3333 0.4754 0.2745 0.3824 0.6580
In-Context-Learning GPT P 0.6667 0.2222 0.2222 0.4595 0.7000 0.5610 0.6957
R 0.4615 0.4000 0.2857 0.6296 0.3684 0.5349 0.7059
F1 0.5455 0.2857 0.2500 0.5312 0.4828 0.5476 0.7007
Table 8: Emotion Recognition Results for Seven Emotion Categories. P: precision, R: recall, F1: F1 score, and Supp: support.
Appendix F Prompt details
F.1 Fine-tuned Llama 2
The general prompt for the Llama chat version is given in Figure 14. The prompts for emotion and cause prediction are given in Fig. 12 and Fig. 13. We provide a specific format for the output so as to ease the post-processing where we extract the first emotion label occurring after the "::" sequence of characters.
Figure 12: Example Prompt for emotion prediction using Llama
Figure 13: Example Prompt for cause prediction using Llama
Figure 14: General Prompt Template for Llama
F.2 ICL-GPT
We devise prompt templates to be used in the LangChain framework. {} represent placeholders to be replaced when making a prompt. Video captioning prompt is given in Fig. 15. Due to rate limits, we had to batch the utterances, thus we may have multiple disjoint descriptions of a conversation. We prompt GPT-3.5 using the prompt in Fig. 16 to stitch the descriptions into a single caption. For explaining the retrieved conversation with emotion annotated, we use the prompt in Fig. 17. The retrieved conversation and explanation are now used as demonstration examples for the emotion recognition prompt in Fig. 18. For an explanation of causes in the retrieved-context window, we use the prompt in Fig. 19. The explanations of the retrieved windows are used as demonstration examples in the prompt for cause prediction within a context window as shown in the prompt in Fig. 20.
Figure 15: Video Captioning Prompt Template
Figure 16: Batched Video Caption Stitching Prompt Template
Figure 17: Emotion Label Explanation Prompt Template
Figure 18: Emotion Recognition with Context Learning Prompt Template
Figure 19: Cause Explanations Prompt Template
Figure 20: Cause Prediction with Context Learning Prompt Template